How many milliliters of a 40% solution of battery acid must be added to 30 milliliters of an 8% solution to make a 20% solution of acid?
Answer by addingup(3677)

0.4x+0.08(30) = 0.2(x+30)
0.4x+2.4 = 0.2x+6 subtract 0.2x and 2.4 on both sides
0.2x = 3.6
x = 3.6/0.2 = 18 ml is how much must be added to make a 20% solution